# Offline Meta-learning for Real-time Bandwidth Estimation

### v2

@@ I Introduction
However, selecting the optimal BWE presents three key challenges: First, QoS and QoE have an ambiguous relationship, e.g., a $50\%$ throughput reduction from $4$ Mbps to $2$ Mbps may not impact user experience due to modern enhancements such as super resolution [6]. Second, network environments fluctuate rapidly, with different BWEs reacting uniquely to transient conditions, complicating fair algorithm comparison [7]. Third, practicality concerns arise as online learning approaches require thousands of live calls [1] with active exploration, potentially degrading QoE by increasing video freeze rates up to $79\%$ [8], making online learning from scratch impractical for live production environments where call quality is paramount.
@@
To address the previously mentioned challenges, we introduce Ivy [^1] A climbing plant known for its ability to thrive in various environments.. Ivy is trained entirely from offline telemetry logs to maximize user QoE based on observed network observations, bridging the gap between QoS and QoE. We leverage Implicit Q-learning (IQL), an offline reinforcement learning (RL) algorithm [9], which allows Ivy to learn from past experiences without the need for live network interactions. Lastly, we design our method to operate at a coarse $6$-second decision interval, enabling Ivy to more effectively monitor the impact of a given BWE on user QoE. In essence, Ivy functions as a metapolicy that dynamically selects the most suitable BWE algorithm based on the observed network conditions, with the goal of maximizing user QoE. We deploy Ivy in Microsoft Teams and utilize a diverse set of traces from fluctuating, burst loss, and stable bandwidth scenarios and nonstationary environments such as Broadband (BB), 5G, and low earth orbit (LEO) networks to analyze the performance of Ivy.
@@ III-B Problem Formulation
RL Formulation. At each decision interval $t$, we construct state $s_{t}=(\vec{d_{t}},\vec{j_{t}},\vec{l_{t}},\vec{u_{t}},\vec{v_{t}},\vec{x_{t}},\vec{z_{t}})$ to capture the network conditions and the previous policy decisions. Building on prior work [2], state $s_{t}$ captures the following normalized QoS signals observed at interval $t$: one-way delay $\vec{d_{t}}$, packet interarrival time $\vec{j_{t}}$, number of lost packets $\vec{l_{t}}$, proportion of audio packets received $\vec{u_{t}}$, proportion of video packets $\vec{v_{t}}$, and receiving rate $\vec{x_{t}}$. These QoS state variables consist of multiple measurements which correspond to a coarse view of the network environment over the previous $6$ seconds, a monitoring window suggested by prior works [7]. Providing the full $6$ seconds of QoS signals at $60$ ms intervals increases the state size by an order of magnitude. To strike a balance between state size and network state granularity, we opt for a more coarse grain approach and collect $10$ measurements, where each sample measurement for each QoS metric represents the average for a $600$ ms window. In contrast to the QoS signals, $\vec{z_{t}}$ captures the previous $5$ metapolicy actions, which enables the policy to capture the longterm impacts of prior policy decisions. We input the $65$ dimensional state $s_{t}$ into our actor network.
@@ III-C Data Collection
Offline learning is bound by the best combination of policies present in the fixed offline dataset. To ensure adequate coverage over the space of metapolicies, we collected logs from $1000$ videoconferencing calls using a random policy. Videoconferencing calls were conducted for $2$ minutes over emulated links where network traces were sampled from diverse scenarios, encompassing fluctuating bandwidth, burst loss patterns, and stable connections with bandwidths ranging from $0.1$ Mbps to $8$ Mbps, as well as real-world nonstationary environments such as BB, 5G, and LEO. To enhance dataset diversity, we uniformly sampled environmental conditions, with packet loss rates ranging from 0 to $0.25$ and round-trip delay times varying from $40$ to $60$ ms.
@@ IV Evaluation
    Ivy effectively handles network nonstationarity, achieving up to $6\%$ video QoE improvements across diverse environments (5G, BB, and LEO) compared to baseline estimators.
@@ IV-A Setup
Environment Setup. We evaluate Ivy on two complementary testbeds with A/B testing. We conduct peer-to-peer videoconferencing calls with BWEs running at 6060ms granularity and metapolicy decisions executing every $6$ seconds. In our primary setup, we utilize a Linux-based network router with netem to emulate various network conditions across randomly selected machine pairs within the same datacenter. Tests run on holdout traces covering fluctuating, burst loss, and stable conditions in both low bandwidth (LBW) and high bandwidth (HBW) environments. For additional validation, we leverage Cloudlab [17], AlphaRTC [18] (an open-source RTC stack), and the Mahimahi network emulation tool [19] and evaluate performance on public network traces [20] from real-world 5G, BB, and LEO satellite environments.
@@
Baselines. We evaluate Ivy against three baseline categories. (1) We compare against the following individual estimators. UKF [2], an unscented Kalman filter optimized for QoS; R3Net [2], an LSTM-based estimator trained via offline RL to maximize video MOS that has previously shown a mix of improvements and regressions across network environments in comparison to production-grade UKF. (2) We compare against online exploration-exploitation and rule-based metapolicies. For exploration-exploitation, we implement $6$-second decision intervals. We follow prior approaches [12] and begin exploration with the more conservative QoS-based BWE UKF. The estimator achieving higher QoS utility during exploration (see Table II for QoS formulations) is selected for the remainder of the call. Additionally, we compare against three rule-based approaches: jitter-based, delta-based, and a combination of the two. The jitter threshold is set to $25$ ms. In contrast, for the delta-based rule, if the difference between UKF and R3Net exceeds one standard deviation from prior estimates, then the policy defaults to the QoS-driven UKF estimator with R3Net running during stable conditions. Rule-based parameters were selected based on empirical observations. (3) We baseline against online RL meta-models trained via Proximal Policy Optimization (PPO) [21], a state-of-the-art online RL algorithm. We train multiple online meta-models on live videoconferencing calls from traces drawn from the same training trace distribution as Ivy.
@@ IV-D Strength in Multiple Estimators
In contrast to our prior evaluations, we now evaluate Ivy in an opensource testbed with AlphaRTC [18]. Using public 5G, BB, and LEO network traces [20] and the Mahimahi emulation tool [19], we collect offline telemetry from $1000$ videoconferencing calls with WebRTC’s Google Congestion Control (GCC) [22], a widely adopted rate control algorithm. Inspired by [8], we then train five new BWEs with offline RL, which we refer to as model-0 through model-$4$. Each BWE is trained on differing subsets of offline telemetry, with no one model being trained on all data. Lastly, we follow the data collection process from Section III and retrain Ivy. We conduct $30$ videoconferencing calls per BWE in this new environment and report our observed QoE values in Figure 4.
Remark: Metalearning enables robust adaption to nonstationary network environments. Rather than relying on a single, generalizable model, Ivy leverages multiple specialized models. Ivy intelligently selects the most suitable BWE to run based on the current network state, leading to a $6\%$ improvement in video QoE on average compared against our baseline estimators in these nonstationary environments. Specifically, we observe no statistically significant regressions in both LEO and 5G cases; however, we observe a $12\%$ improvement in BB. By dynamically selecting the best estimator for the current network conditions, Ivy enhances video QoE in nonstationary network environments.

### v3

@@ Abstract
©2026 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
However, selecting the optimal BWE presents three key challenges: First, QoS and QoE have an ambiguous relationship, e.g., a $50\%$ throughput reduction from 44 Mbps to 22 Mbps may not impact user experience due to modern enhancements such as super resolution [6]. Second, network environments fluctuate rapidly, with different BWEs reacting uniquely to transient conditions, complicating fair algorithm comparison [7]. Third, practicality concerns arise as online learning approaches require thousands of live calls [1] with active exploration, potentially degrading QoE by increasing video freeze rates up to $79\%$ [8], making online learning from scratch impractical for live production environments where call quality is paramount.
@@
To address the previously mentioned challenges, we introduce Ivy [^1] A climbing plant known for its ability to thrive in various environments.. Ivy is trained entirely from offline telemetry logs to maximize user QoE based on observed network observations, bridging the gap between QoS and QoE. We leverage Implicit Q-learning (IQL), an offline reinforcement learning (RL) algorithm [9], which allows Ivy to learn from past experiences without the need for live network interactions. Lastly, we design our method to operate at a coarse 66-second decision interval, enabling Ivy to more effectively monitor the impact of a given BWE on user QoE. In essence, Ivy functions as a metapolicy that dynamically selects the most suitable BWE algorithm based on the observed network conditions, with the goal of maximizing user QoE. We deploy Ivy in Microsoft Teams and utilize a diverse set of traces from fluctuating, burst loss, and stable bandwidth scenarios and nonstationary environments such as Broadband ($B$), 5G, and low earth orbit (LEO) networks to analyze the performance of Ivy.
@@ III-B Problem Formulation
RL Formulation. At each decision interval $t$, we construct state $s_{t}=(\vec{d_{t}},\vec{j_{t}},\vec{l_{t}},\vec{u_{t}},\vec{v_{t}},\vec{x_{t}},\vec{z_{t}})$ to capture the network conditions and the previous policy decisions. Building on prior work [2], state $s_{t}$ captures the following normalized QoS signals observed at interval $t$: one-way delay $\vec{d_{t}}$, packet interarrival time $\vec{j_{t}}$, number of lost packets $\vec{l_{t}}$, proportion of audio packets received $\vec{u_{t}}$, proportion of video packets $\vec{v_{t}}$, and receiving rate $\vec{x_{t}}$. These QoS state variables consist of multiple measurements which correspond to a coarse view of the network environment over the previous 66 seconds, a monitoring window suggested by prior works [7]. Providing the full 66 seconds of QoS signals at $60$ ms intervals increases the state size by an order of magnitude. To strike a balance between state size and network state granularity, we opt for a more coarse grain approach and collect $10$ measurements, where each sample measurement for each QoS metric represents the average for a $600$ ms window. In contrast to the QoS signals, $\vec{z_{t}}$ captures the previous 55 metapolicy actions, which enables the policy to capture the longterm impacts of prior policy decisions. We input the $65$ dimensional state $s_{t}$ into our actor network.
@@ III-C Data Collection
Offline learning is bound by the best combination of policies present in the fixed offline dataset. To ensure adequate coverage over the space of metapolicies, we collected logs from $1000$ videoconferencing calls using a random policy. Videoconferencing calls were conducted for 22 minutes over emulated links where network traces were sampled from diverse scenarios, encompassing fluctuating bandwidth, burst loss patterns, and stable connections with bandwidths ranging from $0.1$ Mbps to 88 Mbps, as well as real-world nonstationary environments such as $B$, 5G, and LEO. To enhance dataset diversity, we uniformly sampled environmental conditions, with packet loss rates ranging from 0 to $0.25$ and round-trip delay times varying from $40$ to $60$ ms.
@@ IV Evaluation
    Ivy effectively handles network nonstationarity, achieving up to $6\%$ video QoE improvements across diverse environments (5G, $B$, and LEO) compared to baseline estimators.
@@ IV-A Setup
Environment Setup. We evaluate Ivy on two complementary testbeds with A/B testing. We conduct peer-to-peer videoconferencing calls with BWEs running at 6060ms granularity and metapolicy decisions executing every 66 seconds. In our primary setup, we utilize a Linux-based network router with netem to emulate various network conditions across randomly selected machine pairs within the same datacenter. Tests run on holdout traces covering fluctuating, burst loss, and stable conditions in both low bandwidth (LBW) and high bandwidth (HBW) environments. For additional validation, we leverage Cloudlab [17], AlphaRTC [18] (an open-source RTC stack), and the Mahimahi network emulation tool [19] and evaluate performance on public network traces [20] from real-world 5G, $B$, and LEO satellite environments.
@@
Baselines. We evaluate Ivy against three baseline categories. (1) We compare against the following individual estimators. UKF [2], an unscented Kalman filter optimized for QoS; R3Net [2], an LSTM-based estimator trained via offline RL to maximize video MOS that has previously shown a mix of improvements and regressions across network environments in comparison to production-grade UKF. (2) We compare against online exploration-exploitation and rule-based metapolicies. For exploration-exploitation, we implement 66-second decision intervals. We follow prior approaches [12] and begin exploration with the more conservative QoS-based BWE UKF. The estimator achieving higher QoS utility during exploration (see Table II for QoS formulations) is selected for the remainder of the call. Additionally, we compare against three rule-based approaches: jitter-based, delta-based, and a combination of the two. The jitter threshold is set to $25$ ms. In contrast, for the delta-based rule, if the difference between UKF and R3Net exceeds one standard deviation from prior estimates, then the policy defaults to the QoS-driven UKF estimator with R3Net running during stable conditions. Rule-based parameters were selected based on empirical observations. (3) We baseline against online RL meta-models trained via Proximal Policy Optimization (PPO) [21], a state-of-the-art online RL algorithm. We train multiple online meta-models on live videoconferencing calls from traces drawn from the same training trace distribution as Ivy.
@@ IV-D Strength in Multiple Estimators
In contrast to our prior evaluations, we now evaluate Ivy in an opensource testbed with AlphaRTC [18]. Using public 5G, $B$, and LEO network traces [20] and the Mahimahi emulation tool [19], we collect offline telemetry from $1000$ videoconferencing calls with WebRTC’s Google Congestion Control (GCC) [22], a widely adopted rate control algorithm. Inspired by [8], we then train five new BWEs with offline RL, which we refer to as model-0 through model-44. Each BWE is trained on differing subsets of offline telemetry, with no one model being trained on all data. Lastly, we follow the data collection process from Section III and retrain Ivy. We conduct $30$ videoconferencing calls per BWE in this new environment and report our observed QoE values in Figure 4.
Remark: Metalearning enables robust adaption to nonstationary network environments. Rather than relying on a single, generalizable model, Ivy leverages multiple specialized models. Ivy intelligently selects the most suitable BWE to run based on the current network state, leading to a $6\%$ improvement in video QoE on average compared against our baseline estimators in these nonstationary environments. Specifically, we observe no statistically significant regressions in both LEO and 5G cases; however, we observe a $12\%$ improvement in $B$. By dynamically selecting the best estimator for the current network conditions, Ivy enhances video QoE in nonstationary network environments.